### v2

@@ Abstract
active learning, object detection, reinforcement learning, policy gradient.
@@ I Introduction
Figure 1: Training and selection pipelines of Mean-AP Guided Reinforced Active Learning (MGRAL) for object detection. (a) Data sampling agent training phase. An RL-based agent is trained using $\Delta$mAP as reward to prioritize data batch that improves detection neural network performance most. (b) Data selection with trained sampling agent. The trained agent is used as the query strategy to select the most informative data batch for annotation and detection model fine-tuning.
@@
To estimate the potential change on mAP when including unlabeled samples, MGRAL employs an unsupervised model. This approach is effective as it captures global trends sufficient for identifying potentially impactful data. The unsupervised model’s output need not be absolutely correct labels; rather, it aims to correctly estimate which data significantly differ from the model’s current decision boundaries or distributions. This relative information is crucial for guiding the active learning process in selecting the most informative samples for labeling. Additionally, to address the extensive training time required by the reinforcement learning agent, which necessitates retraining the semi-supervised model iteratively, we implement fast lookup tables for acceleration. Empirical results demonstrate our method’s efficiency on both Pascal VOC [7] and MS COCO [8] datasets.
@@ III-B Overview of MGRAL Pipeline
Our proposed method MGRAL derives its query strategy by utilizing mean Average Precision of the task learner to maximize the impact of selected batches.
As shown in Figure 1, MGRAL integrates a reinforcement learning-based sampling agent (controller) into the typical pool-based active learning pipeline. This agent addresses a fundamental challenge in active learning for object detection: the difficulty in directly optimizing batch selection based on mAP improvement. In conventional approaches, it’s challenging to establish a direct link between the selected batch of samples and the resulting mAP improvement due to the discrete and combinatorial nature of batch selection. The reinforcement learning agent overcomes this by learning to navigate this complex, discrete search space, using mAP improvements as feedback. This approach allows the agent to effectively learn the correlation between batch selections and their impact on mAP, without requiring a differentiable relationship.
Figure 2: Data sampling agent architecture.
@@ III-C MGRAL Data Sampling Agent
The MGRAL data sampling agent employs a Long Short-Term Memory (LSTM) [28] network to optimize the selection of informative images from the unlabeled pool. As shown in Figure 2, each image is first processed through a feature extraction network $\Phi$, which derives a feature vector using a pre-trained detection model.
The LSTM Modules operate in a parameter-sharing mode across all images, enhancing scalability and preventing gradient vanishing. Moreover, we compute the embedding vector for each image by concatenating the feature vector with the decision vector from the preceding unit. This approach integrates both new data and historical decisions. The interaction between the LSTM Module outputs and the decision-making process is depicted as:
@@
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation operation, $\Psi_{i}$ is the decoding network for the $i$-th image. This structure ensures each image’s selection considers its potential contribution to the model’s performance, as quantified by changes in mAP.
@@ III-D Performance-Driven Reward Design
The core of our approach lies in directly associating data sampling decisions with the performance improvements of the detection model. We define the agent’s reward as the variation in mean Average Precision ($\Delta\text{mAP}$). For each training iteration $i$, we combine the current labeled dataset $X_{L}$ with the newly selected unlabeled samples $X_{S}^{i}$. A new detector is then trained on this combined set with $\text{mAP}_{i}$ computed. The reward is calculated as $\Delta\text{mAP}=\text{mAP}_{i}-\text{mAP}_{i-1}$.
Given that newly selected samples are initially unlabeled, we employ a semi-supervised detection model to estimate the detector’s performance. This proxy model facilitates the estimation of mAP in the absence of fully labeled data, crucial for our active learning setting. The semi-supervised model captures global trends of unlabeled data distribution sufficient for identifying potentially impactful data, providing relative information about data informativeness rather than accurate predictions or pseudo-labels. Importantly, even if the semi-supervised model’s performance is not optimal, it does not significantly impact our method’s effectiveness as we only require a reasonable ranking of data informativeness. While semi-supervised models cannot completely replace supervised learning, this active learning approach allows us to efficiently select valuable data for annotation, optimizing the trade-off between labeling effort and model performance.
@@ III-E Training Process and Optimization
The training process of MGRAL involves iterative refinement of the data sampling agent. In each iteration, the agent first selects a candidate set from the unlabeled pool, which is then used to update the semi-supervised detection model. The model’s performance, measured by estimated mAP, guides the agent’s learning. To stabilize training, we employ a moving average mAP (mAPref) as a baseline:
@@
The agent’s loss is calculated as the negative difference between the current mAP and this reference:
@@
This process repeats for $max\_search\_iters$ times, progressively improving the agent’s sample selection strategy.
@@ III-F Acceleration Technique
MGRAL requires frequent retraining of the semi-supervised detector to estimate mAP for each potential batch selection, leading to significant computational costs. To mitigate this extensive training duration, we employ a fast lookup table technique. We pre-compute and store model performances for a series of labeled dataset sizes, incrementing by the batch size used in each active learning cycle, from the initial set up to budget $B$. During training, instead of retraining the semi-supervised model for each new batch selection, we estimate mAP by comparing the visual features of the selected batch to those in the lookup table.
Specifically, we use L2 distance (outperform cosine similarity) to measure similarity between the visual representations of the selected batch and the pre-computed datasets. The mAP is approximated through a weighted summation of the most similar pre-computed results, with weights inversely proportional to the measured distances to account for confidence. If all similar pre-computed results exceed a set distance threshold (the mean feature distance minus one standard deviation), we revert to training the semi-supervised model directly for an accurate mAP. This approach significantly accelerates the training process while maintaining accuracy when needed.
@@
TABLE II: Overall time cost comparison. First active learning cycle on VOC.
Method Training Time Inference Time
Entropy 0 5 min
MIAL [3] 7 h 13 min 43 min
EBAL [5] 0 181 min
Ours (9 h) + 7 h 8 min 0.5 min
@@ IV-A Experiment Details
Detector and Semi-supervise Settings: On PASCAL VOC, we use SSD [29] detector with ISD-SSD [30], training for 300 epochs (learning rate: 0.001, reduced to 0.0001 after 240 epochs; batch size: 32). For COCO, we employ RetinaNet [31] detector with SED-SSOD [32] for 180k iterations (batch size: 8; learning rate: 0.02, dropped by 0.1 at the 120k and 160k).
Data Sampling Agent Configuration: The controller uses a 257-dimensional vector (256 for image embeddings, 1 for previous selection score) with an LSTM of the same size. We optimize using Adam [33] (learning rate: $3.5\times 10^{-4}$).
RL Training Settings: The maximum search iterations for each cycle are set to 2,200 for Pascal VOC and 800 for MS COCO, with a baseline decay factor of 0.5 for mAPref.
@@
IV-B Overall Performance
We compare MGRAL with various baselines including random sampling, entropy sampling, Core-set [14], CDAL [2], LL4AL [1], MIAL [3, 27], and the state-of-the-art EBAL [5]. Experiments were conducted on GTX 1080Ti GPUs for PASCAL VOC and Tesla V100 GPUs for MS COCO.
As shown in Fig.3, MGRAL consistently outperforms all methods on PASCAL VOC, demonstrating the effectiveness of our mAP-guided approach. On MS COCO, MGRAL beats all other baselines across all cycles except EBAL [5]. MGRAL exhibits the steepest performance curve, ultimately surpassing EBAL in the final cycle. This curve indicates the mAP-guided strategy and direct mAP optimization by policy gradient offer unique advantages in active learning for object detection.
Figure 4: Selected samples of first cycle on Pascal VOC.
IV-C Qualitative Analysis
@@
IV-D Efficiency and Ablation Analysis
The fast lookup table significantly reduces training time from 4800 to 3 minutes for ten iterations on four GTX 1080Ti GPUs (Tab.3). Compared to entropy sampling, MIAL [3], and EBAL [5], MGRAL shows competitive overall time efficiency on VOC, offering shorter inference times in subsequent cycles despite a longer initial setup (Tab.3). These demonstrate MGRAL’s long-term efficiency in active learning, effectively balancing setup costs with improved subsequent performance.
@@ V Conclusion
We presented MGRAL, a Mean-AP Guided Reinforced Active Learning for object detection. By using $\Delta\text{mAP}$ as reward, our method aligns batch selection with performance improvement, addressing the non-differentiable nature of this process. MGRAL demonstrates strong results on VOC and COCO. While our fast lookup table effectively accelerates training, it remains a preliminary solution open for further optimization. This work establishes a new paradigm in active learning for object detection, demonstrating the effectiveness of performance-driven reinforcement learning in this area.

### v3

@@ Abstract
active learning, reinforcement learning, data mining, object detection.
@@ I Introduction
Figure 1: Training and selection pipelines of Mean-AP Guided Reinforced Active Learning (MGRAL) for object detection. (a) Data sampling agent training phase: The RL-based agent learns to select informative samples using $\Delta$mAP as reward, where performance gains are efficiently estimated through a semi-supervised detector. (b) Active selection phase: The trained agent directly processes the unlabeled pool to select samples for oracle annotation, which are then added to the labeled pool for detector improvement.
@@
To estimate the potential change on mAP when not having labels of selected samples, MGRAL employs unsupervised model to approximate. This approach is effective as it captures global trends sufficient for identifying potentially impactful data. The unsupervised model’s output need not be absolutely correct labels; rather, it aims to correctly estimate which data significantly differ from the model’s current decision boundaries or distributions. This relative information is crucial for guiding the active learning process in selecting the most informative samples for labeling. Additionally, to address the extensive training time required by the reinforcement learning agent, which necessitates retraining the semi-supervised model iteratively, we implement fast lookup tables for acceleration. Empirical results demonstrate our method’s efficiency on both Pascal VOC [7] and MS COCO [8] datasets.
@@ III-B Overview of MGRAL Pipeline
Our proposed method MGRAL introduces a novel approach to active learning for object detection that directly optimizes mean Average Precision (mAP) through reinforcement learning. The core innovation lies in integrating a reinforcement learning-based sampling agent into the conventional pool-based active learning pipeline, which learns to select the most informative samples by directly optimizing for mAP improvement.
As shown in Figure 1, during training, MGRAL operates as a nested optimization process. The outer loop trains the sampling agent iteratively, while the inner loop focuses on calculating mAP improvement. Specifically, for each active learning round, (Inner Loop) we first construct a lookup table by parallel training multiple detector variants (Section III-F). This table serves as an efficient approximation mechanism for estimating performance gains. (Outer Loop) The sampling agent, detailed in Section III-C, then undergoes training iterations where it selects candidate sets from the unlabeled pool and receives feedback based on the estimated mAP improvements (Section III-D). This feedback drives the policy gradient updates of the sampling agent through our stabilized optimization process (Section III-E).
During inference (active sample selection), the trained sampling agent directly processes the unlabeled pool through its LSTM-based architecture. For each image, it computes a selection score based on both the image’s features and the historical selection context. The system then selects the top-scoring images within the specified budget for human annotation.
This design allows MGRAL to effectively bridge the gap between sample selection and model performance improvement, while maintaining computational efficiency through our lookup table acceleration technique. The following sections detail each component of this framework and their interactions.
Figure 2: Data sampling agent architecture. The agent processes unlabeled images sequentially through three main components: (1) an Image Feature Encoder that extracts visual representations from input images, (2) a series of parameter-shared LSTM modules that process image embeddings while maintaining temporal dependencies through hidden states ($h_{i}$) and cell states ($c_{i}$), and (3) individual decoder networks that output selection scores for each image. A dummy head is used to initialize the first LSTM state. Each image embedding combines the feature representation with the previous decision, enabling considering both visual content and selection history.
@@ III-C MGRAL Data Sampling Agent
The sampling agent adopts a Neural Architecture Search (NAS) inspired design, utilizing a Long Short-Term Memory (LSTM) [28] network for sequential sample selection. As shown in Figure 2, each image is first processed through a feature extraction network $\Phi$, which derives a feature vector using a pre-trained detection model. Then the extracted features then flow through parameter-shared LSTM modules that ensure scalability while preventing gradient vanishing. The final component is a 2-layer MLP decoder that outputs selection scores for each sample.
The agent’s architecture integrates both current features and historical decisions through combining each image’s embedding vector with the decision vector from the preceding unit, enabling the agent to consider both new data characteristics and previous selection patterns. This interaction is formally expressed as:
@@
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation operation, $\Psi_{i}$ is the decoding network for the $i$-th image. During operation, the agent uses these computed scores to select the top-$B$ samples from the unlabeled pool, where $B$ is the active learning budget. This selection mechanism forms the basis for the reward computation and policy gradient updates detailed in Sec. III-E.
@@ III-D Performance-Driven Reward Design
The reward mechanism forms the cornerstone of our approach by directly linking sample selection to detector performance through mAP improvement ($\Delta\text{mAP}$). During each training iteration i, the system combines the current labeled dataset $X_{L}$ with newly selected samples $X_{S}^{i}$. A new detector is trained on this combined set to compute $\text{mAP}_{i}$, with the reward calculated as $\Delta\text{mAP}=\text{mAP}_{i}-\text{mAP}_{i-1}$.
To address the challenge of evaluating unlabeled samples during the selection process (sampling agent training & inference), we employ a semi-supervised detection model as a proxy for performance estimation. Specifically, during active selection of unlabeled data, we need to predict the performance gain that would result from oracle labeling, without having access to the actual annotations. We achieve this by using a semi-supervised model trained on $X_{i-1}+\text{(unlabeled)}X_{B}$ to approximate the performance gain of a supervised model trained on $X_{i-1}+\text{(labeled)}X_{B}$, where $X_{B}$ represents the candidate unlabeled batch.
Importantly, we utilize the semi-supervised model solely as a mechanism to capture global trends in the unlabeled data distribution and provide relative information about data informativeness. This approach enables us to rank and identify potentially impactful samples without requiring perfect predictions or pseudo-labels, making it an effective proxy for estimating the potential value of unlabeled samples once they are properly annotated.
@@ III-E Training Process and Optimization
The training process implements a robust optimization strategy centered on iterative refinement of the data sampling agent. For each iteration, the agent produces selection scores for unlabeled samples through its LSTM architecture and selects the top-$B$ scoring samples to form a binary mask. Using policy gradient, we enable reward gradients to flow back through this discrete selection process by multiplying the mask with the downstream reward gain, allowing effective update of the agent’s parameters.
To stabilize the policy gradient training, we employ a moving average baseline mechanism that adapts to the changing performance landscape. Let $\text{mAP}_{i}$ denote the mean Average Precision achieved by the detector after incorporating the i-th batch of selected samples. We maintain a reference baseline $\text{mAP}_{ref}$ that tracks the average performance trend:
@@
where $\lambda$ is a momentum coefficient that controls the update rate of the reference baseline. This dynamic baseline helps normalize the reward signal and reduce variance in the policy gradient updates.
The agent’s loss is then defined relative to this adaptive baseline:
@@
Here, $\text{mAP}_{i}$ represents the estimated performance after incorporating the current batch selection, while $\text{mAP}_{ref}$ serves as a moving average of historical performance values. By subtracting the baseline from the current performance, we obtain a more stable learning signal that indicates whether the current batch selection is better or worse than the recent average. This normalized reward, combined with the policy gradient approach, enables effective optimization of the sampling agent’s parameters through the discrete selection process.
The training process continues for a fixed number of iterations or until convergence, with the agent progressively learning to identify more informative samples based on the relative performance improvements they provide. This policy gradient optimization framework allows the agent to learn effective sample selection strategies despite the discrete nature of the selection process and the non-differentiable mAP metric.
@@ III-F Acceleration Technique
A critical challenge in our approach lies in the computational cost of the inner loop optimization. Without acceleration, each iteration of the sampling agent training would require a complete retraining of the semi-supervised detector to estimate mAP for every potential batch selection. This would result in an prohibitive computational overhead of 2200 (VOC) or 800 (COCO) sequential detector training iterations for each round of active learning.
To address this challenge, we introduce an efficient lookup table-based acceleration strategy. The system constructs this table by training $M$ semi-supervised models in parallel, each utilizing the current labeled data ($X_{labeled}$) plus a randomly selected batch of unlabeled data ($\widetilde{X}_{B}$) within the active learning budget $B$. This parallel pre-computation serves as a basis for rapid performance estimation during the sampling agent’s training iterations.
During the agent training, instead of retraining the detector, performance estimation leverages the Wasserstein distance [29] (one of optimal transport distance) to measure similarity between the visual representations of selected batches and pre-computed datasets. The system computes weights inversely proportional to measured distances, enabling accurate mAP approximation through weighted summation of the most similar pre-computed results. In cases where similarity measures exceed a predetermined threshold (mean feature distance minus one standard deviation), the system falls back to direct training for precise mAP calculation.
This acceleration approach dramatically reduces computational overhead, achieving a 1000×1000$\times$1000 × speedup by transforming the requirement from thousands of sequential detector training iterations to a single parallel training phase per round of data selection. This efficiency gain makes the proposed method practically feasible while maintaining its effectiveness in selecting informative samples.
@@
TABLE II: Efficiency analysis. All results are about models on PASCAL VOC.
Method Training Time of One Cycle Inference Time of One Cycle # of params
Entropy 0 5 min 26.3 M
MIAL [3] 7 h 13 min 43 min 31.9 M
EBAL [5] 0 181 min 26.5 M
Ours (9 h) + 7 h 8 min 0.5 min 33.0 M
Figure 4: Visualizations of most representative selected samples of different methods during first cycle on Pascal VOC.
@@ IV-A Experiment Details
Data Sampling Agent Settings (outer loop): The controller uses a 257-dimensional vector (256 for image embeddings, 1 for previous selection score) with an LSTM of the same size. We optimize using Adam [30] (learning rate: $3.5\times 10^{-4}$). The maximum search iterations for each cycle are set to 2,200 for Pascal VOC and 800 for MS COCO, with a baseline decay factor of 0.5 for mAPref.
Detector Settings (inner loop): On PASCAL VOC, we use SSD [31] detector with ISD-SSD [32], training for 300 epochs (learning rate: 0.001, reduced to 0.0001 after 240 epochs; batch size: 32). For COCO, we employ RetinaNet [33] detector with SED-SSOD [34] for 180k iterations (batch size: 8; learning rate: 0.02, dropped by 0.1 at the 120k and 160k).
@@
IV-B Quantitative Performance
We compare MGRAL with various baselines including random sampling, entropy sampling, Core-set [14], CDAL [2], LL4AL [1], MIAL [3, 27], and EBAL [5]. Experiments were conducted on GTX 1080Ti GPUs for PASCAL VOC and Tesla V100 GPUs for MS COCO.
As shown in Fig.3, MGRAL consistently outperforms all methods on PASCAL VOC, demonstrating the effectiveness of our mAP-guided approach. The superior performance can be attributed to our direct optimization of mAP through policy gradient, which naturally aligns the selection strategy with the task’s evaluation metric. On MS COCO, MGRAL shows competitive performance, achieving the steepest performance curve and ultimately surpassing EBAL in the final cycle, though it lags behind EBAL in early cycles. This pattern suggests that while our method excels at fine-grained selection, it may benefit from hierarchical sampling strategies when dealing with large-scale unlabeled pools, potentially using clustering to reduce the search space before applying our selection mechanism.
Our approach’s effectiveness stems from two key aspects: first, using mAP as the reward signal provides direct optimization of the detection task’s primary metric; second, the reinforcement learning framework enables flexible adaptation to any downstream task by simply changing the reward metric, making our method potentially generalizable beyond detection.
IV-C Ablation on Visualization Analysis
@@
IV-D Ablation on Efficiency Analysis
For time complexity, as shown in Tab.3, the fast lookup table significantly reduces training time from 4800 to 3 minutes for ten iterations on four GTX 1080Ti GPUs. While this initial training requires more setup time than baselines like entropy sampling, MIAL [3], and EBAL [5], MGRAL demonstrates superior efficiency in subsequent cycles through shorter inference times (Tab.3). Importantly, this training overhead is acceptable for real-world deployment scenarios, where data collection, selection, and labeling typically occur in offline cycles rather than requiring real-time responses.
In terms of space complexity, our method maintains a competitive memory footprint, with the model requiring $33.0$M parameters compared to EBAL’s $26.5$M and MIAL’s $31.9$M parameters (Tab.3). The additional lookup table introduces minimal overhead, requiring only $12$KB of storage while enabling substantial computational acceleration.
TABLE III: Ablation on the use of lookup table on VOC.
Table Size 1000 2000 3000 4000 5000 6000 7000 8000 9000 10000
50 records 56.41 62.95 67.22 70.33 72.12 73.04 74.04 74.83 75.57 75.82
100 records 56.41 62.91 67.56 69.82 72.38 72.82 73.90 74.46 75.02 75.57
200 records 56.41 63.26 68.09 70.30 72.05 73.16 73.96 74.94 75.41 75.85
IV-E Ablation on the Use of Lookup Table
To investigate the impact and robustness of our lookup table acceleration technique, we conduct experiments with varying numbers of pre-computed records in the fast lookup table. As shown in Tab.III, we evaluate the performance using 50, 100, and 200 records across different training stages (1k-10k images) on VOC dataset.
The results demonstrate that as long as the lookup table maintains a reasonable size, our method can achieve superior performance. While 200 records yields slightly better results (75.85 mAP at 10k images), even with just 50 records, the method achieves strong performance (75.82 mAP). This suggests that our acceleration strategy is effective with a modest number of pre-computed records, offering substantial speedup while maintaining performance advantages.
@@ V Conclusion
We presented MGRAL, a Mean-AP Guided Reinforced Active Learning framework for object detection. By using $\Delta\text{mAP}$ as reward, our method effectively aligns batch selection with detection performance improvement, addressing the non-differentiable nature of this process through policy gradient optimization. The integration of semi-supervised detection and lookup table-based acceleration enables practical deployment while maintaining efficiency. MGRAL achieves consistent improvements over prior methods on PASCAL VOC and demonstrates competitive performance on MS COCO, establishing a new paradigm that combines performance-driven reinforcement learning with efficient active learning for object detection.
Future direction: Explore early stopping techniques for faster mAP estimation and replacing lookup tables with prediction networks for online reinforcement learning.